# Constraining the position of the CEP through the Speed of Sound in the LSMq

### v2

@@ Abstract
One of such examples, the Linear Sigma Model with quarks (LSMq), has been successfully used to reproduce the chiral transition of strongly interacting matter under extreme conditions by appropriately taking into account the plasma screening effects in the dynamics near the transition [25, 26, 27, 28, 29, 30, 31]. The behavior of the effective potential including ring diagrams [32] with respect to order parameter after spontaneous symmetry breaking is often employed to locate the CEP in thermodynamical parameter space. A favorite strategy to this end is to derive the statistical properties of the system from the fluctuations of the order parameter with respect to parameters such as energy of the collision, temperature and baryon chemical potential [30]. With the aim of extending these analysis, in this letter we focus on the behavior of transport quantities derived from the thermodynamics of the system near the transition. In particular, we consider the speed of sound for homentropic, $c_{s}$, isobaric $c_{\rho}$, and isentropico $c_{s/\rho}$ processes. It is expected that exactly at the CEP, $c_{s/\rho}$ vanishes and that would be the smoking gun to locate this point in the phase diagram. To set constraints on the values of the couplings of the LSMq, we demand that $c_{s/\rho}\simeq c_{\rho}$ approaches the lattice extracted value at $\mu_{B}=0$ [33] and see how far it departures from zero as we permit the temperature and chemical potential to evolve in the phase diagram.
@@
where $\sigma$ represents the sigma boson, $\vec{\pi}$ is the pion triplet, with the neutral pion $\pi^{0}=\pi_{3}$ and the charged pions $\pi^{\pm}=(\pi_{1}\mp\pi_{2})/\sqrt{2}$, and $\psi$ is an isospin $SU(2)$ doublet representing the $u$ and $d$ quarks $\vec{\tau}$ are the Pauli matrices in isospin space. Mesons in this model are coupled with a four body contact interaction with strength $\lambda$ and with a three-point vertex with quarks with coupling strength $g$. Moreover, $a$ is a dimensionful parameter with mass units 1. Notice that when quarks remain massless, the Lagrangian (1) is invariant under chiral transformations. Its effective potential at high temperature up to order ring reads
@@
In Fig. 4 we depict the behavior of $c_{s/\rho}^{2}$ for different critical values $\mu_{B}^{c}$ and the corresponding critical temperatures $T_{c}$. Again, different curves are for the parameters of $N_{f}=2$ and $N_{f}=2+1$ in LQCD. For $N_{f}=2$, the CEP can be located and a sharp fall-off of $c_{s/\rho}^{2}$ is noticeable which can be used as a signal to dectect its position. This is not the case of $N_{f}=2+1$, where no evidence of first order transition can be found because the hight-$T$ approximations breaks before reaching the CEP.
In this letter we re-examined the LSMq in connection with the physics of the QCD phase diagram and the existence of the CEP. We have put forward a strategy to fix the couplings $g$ and $\lambda$ as well as the parameter $a$ in the Lagrangian (1) from three requirements: (i) the condition of continuous phase transitions at high temperature, (5), (ii) the curvature of the critical line in the phase diagram according to LQCD, (6), and (iii) demanding that the speed of sound $c_{s/\rho}^{2}$ corresponds to the measured value. We consider that such a strategy allows to further constrain the free parameter space and discard some approximations in the need for more educated analysis of the phase diagram through effective models of this kind. As compared with a previous analysis by some of us regarding the collision energy dependence in the position of the CEP [30], we have narrowed down the range of values of $\lambda$ to a single point. We stress that the attainable values of $c_{s/\rho}$ in that work are significantly separated from the measured value. In this regard, our demand that the parameter choice renders the measured value of the speed of sound is in fact a tight constraint that any treatment of the effective potential of the LSMq should fulfill.

### v3

@@ Abstract
One of such examples, the Linear Sigma Model with quarks (LSMq), has been successfully used to reproduce the chiral transition of strongly interacting matter under extreme conditions by appropriately taking into account the plasma screening effects in the dynamics near the transition [25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35]. The behavior of the effective potential including ring diagrams [36] with respect to order parameter after spontaneous symmetry breaking is often employed to locate the CEP in thermodynamical parameter space. A favorite strategy to this end is to derive the statistical properties of the system from the fluctuations of the order parameter with respect to parameters such as energy of the collision, temperature and baryon chemical potential [33]. Numerous studies have employed the linear sigma model or other effective models to analyze various thermal properties of the medium, such as the speed of sound, bulk viscosity, transport coefficients, among others [37, 38, 39, 40, 41, 42]. In this letter, we focus on the behavior of transport quantities derived from the thermodynamics of the system near the transition. In particular, we consider the speed of sound for homentropic, $c_{s}$, isobaric $c_{\rho}$, and isentropico $c_{s/\rho}$ processes. It is expected that exactly at the CEP, $c_{s/\rho}$ vanishes and that would be the smoking gun to locate this point in the phase diagram. To set constraints on the values of the couplings of the LSMq, we demand that $c_{s/\rho}\simeq c_{\rho}$ approaches the lattice extracted value at $\mu_{B}=0$ [43] and see how far it departures from zero as we permit the temperature and chemical potential to evolve in the phase diagram.
@@
where $\sigma$ represents the sigma boson, $\vec{\pi}$ is the pion triplet, with the neutral pion $\pi^{0}=\pi_{3}$ and the charged pions $\pi^{\pm}=(\pi_{1}\mp\pi_{2})/\sqrt{2}$, and $\psi$ is an isospin $SU(2)$ doublet representing the $u$ and $d$ quarks $\vec{\tau}$ are the Pauli matrices in isospin space. Mesons in this model are coupled with a four body contact interaction with strength $\lambda$ and with a three-point vertex with quarks with coupling strength $g$. Moreover, $a$ is a dimensionful parameter with mass units 1. Notice that when quarks remain massless, the Lagrangian (1) is invariant under chiral transformations. This Lagrangian provides an effective description of the QCD phase diagram in terms of the chiral phase transition. Working in the strict chiral limit, to allow for a spontaneous symmetry breaking, we let the $\sigma$ field to develop a vacuum expectation value $v$, namely $\sigma\rightarrow\sigma+v$, this vacuum expectation value is identified as the order parameter of the theory. After this shift, the Lagrangian can be rewritten as
${\cal L}$ $=$ $\frac{1}{2}(\partial_{\mu}\sigma)^{2}+\frac{1}{2}(\partial_{\mu}\vec{\pi})^{2}-\frac{1}{2}\left(3\lambda v^{2}-a^{2}\right)\sigma^{2}$ (2)
$-\frac{1}{2}\left(\lambda v^{2}-a^{2}\right)\pi^{2}+\frac{a^{2}}{2}v^{2}-\frac{\lambda}{4}v^{4}$
$+i\bar{\psi}\gamma^{\mu}\partial_{\mu}\psi+gv\bar{\psi}\psi+{\cal L}_{I}^{b}+{\cal L}_{I}^{f},$
in terms of ${\cal L}_{I}^{b}$ and ${\cal L}_{I}^{f}$, which are, respectively, given by
${\cal L}_{I}^{b}$ $=$ $-\frac{1}{4}\left[\left(\sigma^{2}+(\pi^{0})^{2}\right)^{2}+4\pi^{+}\pi^{-}\left(\sigma^{2}+(\pi^{0})^{2}+\pi^{+}\pi^{-}\right)\right]$
${\cal L}_{I}^{f}$ $=$ $-g\bar{\psi}\left(\sigma+i\gamma_{5}\cdot\mathbf{\pi}\right)$ (3)
where the terms in Eq. (3) describe the interactions among the fields $\sigma$, $\mathbf{\pi}$ and $\psi$, after symmetry breaking. From Eq. (2) notice that the $\sigma$, the three pions and the quarks have masses given by
$m_{\sigma}^{2}$ $=$ $3\lambda v^{2}-a^{2}$
$m_{\pi}^{2}$ $=$ $\lambda v^{2}-a^{2}$
$m_{f}$ $=$ $gv$ (4)
Note that in Eq. (4), the squares of the tree-level bosons masses can either disappear or even turn negative as the continuous order parameter $v$ traverses its range. The physical masses, comprising their thermal components, are determined when $v$ adopts the value obtained from minimizing the effective potential. While our focus in this study has not been explicitly directed towards elucidating the behavior of these thermal masses, our findings align seamlessly with those of prior research, such as the study referenced in [44], when accounting for the thermal contributions to these masses through their self-energies (refer to Eq. (6) below), albeit in our case, under the high temperature regime.
The effective potential associated to (2) at high temperature, up to order ring, reads
@@
In Fig. 4 we depict the behavior of $c_{s/\rho}^{2}$ for different critical values $\mu_{B}^{c}$ and the corresponding critical temperatures $T_{c}$. Again, different curves are for the parameters of $N_{f}=2$ and $N_{f}=2+1$ in LQCD. For $N_{f}=2$, the CEP can be located and a sharp fall-off of $c_{s/\rho}^{2}$ is noticeable which can be used as a signal to detect its position. This is not the case of $N_{f}=2+1$, where no evidence of first order transition can be found because the high-$T$ approximations breaks before reaching the CEP.
In this letter we re-examined the LSMq in connection with the physics of the QCD phase diagram and the existence of the CEP. Although it is often desirable to include the effects of strangeness and nuclear matter and their role in the location of the CEP, it is well established that neither one of these have a dominant thermodynamic contribution near criticality. Thus, we decided simply omit them from our considerations. Rather, we have put forward a strategy to fix the couplings $g$ and $\lambda$ as well as the parameter $a$ in the Lagrangian (1) from three requirements: (i) the condition of continuous phase transitions at high temperature, (8), (ii) the curvature of the critical line in the phase diagram according to LQCD, (9), and (iii) demanding that the speed of sound $c_{s/\rho}^{2}$ corresponds to the measured value. We consider that such a strategy allows to further constrain the free parameter space and discard some approximations in the need for more educated analysis of the phase diagram through effective models of this kind. As compared with a previous analysis by some of us regarding the collision energy dependence in the position of the CEP [33], we have narrowed down the range of values of $\lambda$ to a single point. We stress that the attainable values of $c_{s/\rho}$ in that work are significantly separated from the measured value. In this regard, our demand that the parameter choice renders the measured value of the speed of sound is in fact a tight constraint that any treatment of the effective potential of the LSMq should fulfill.